OpenStudy (anonymous):

If 7(f-1)=45,then what does f=?

OpenStudy (anonymous):

52/7

jimthompson5910 (jim_thompson5910):

\[\large 7(f-1)=45\] \[\large 7f-7=45\] \[\large 7f=45+7\] \[\large 7f=52\] \[\large f=\frac{52}{7}\]